If a*b=3a-6b+2, find (5*2)*7
Answer by Edwin McCravy(20054)

a*b = 3*a - 6*b + 2, find (5*2)*7

First find (5*2) by substituting a=5 and b=2 in

a*b = 3*a - 6*b + 2
5*2 = 3*5 - 6*2 + 2 
5*2 =  15 -  12 + 2
5*2 =         3 + 2
5*2 =           5

Since 5*2 = 5,

     (5*2)*7  becomes 
         5*7

So we find 5*7 by substituting a=5 and b=7 in

a*b = 3*a - 6*b + 2
5*7 = 3*5 - 6*7 + 2 
5*7 =  15 -  42 + 2
5*7 =       -27 + 2
5*7 =         -25

Edwin